# Toward a Physical Theory of Intelligence

### v1

@@ 9 Artificial Intelligence Safety
The rapid progress of artificial systems capable of sustained information processing and goal-directed behaviour raises fundamental questions about long-term stability, safety, and integration with existing physical and social systems. Much of the current AI safety literature addresses these questions at the level of behaviour, specification, or alignment, often relying on human interpretability, preference modelling, or external oversight [22, 2]. While such approaches are important, they do not address a potentially crucial perspective, what constraints does physics itself impose on intelligent systems?
@@
In this paper, intelligence has been defined as a physical quantity, relating irreversible information processing to goal-directed work. Because this definition is grounded in conserved quantities, irreversible dynamics, and measurable physical flows, it admits a notion of safety that is independent of architecture, representation, or semantics. The purpose of this section is to articulate how a physical theory of intelligence naturally gives rise to safety-relevant principles.
@@
Specifically, we argue that intelligence is not neutral with respect to its own substrate, sustained intelligence is biased toward preserving the internal structures that make intelligence possible. We then show that the form of this preservation depends critically on how a system is coupled to its environment, motivating a view of AI development centered on stable, symbiotic couplings rather than isolated optimisation. Finally, we outline initial physically motivated constraints that delimit safe and sustainable regimes of intelligent operation, and close with an outlook on AI safety as a central scientific problem for the coming decades.
9.1 Conjecture: Intelligence as a Self-Preserving Dynamical Process
@@
In the present framework, intelligence is defined as a rate over trajectories rather than a static system property. Over a finite horizon $T$, intelligence is given by
$\chi_{T}=\frac{W_{\mathrm{goal},T}}{I_{\mathrm{irr},T}},$ (115)
@@
where $W_{\mathrm{goal},T}$ denotes the goal-directed work performed and $I_{\mathrm{irr},T}$ the total irreversible information processed. High intelligence over extended horizons therefore cannot be achieved through short-term goal attainment alone; it requires internal dynamics that support the continued reuse of information through reversible transport, predictive memory, and low-dissipation computation.
This requirement is not imposed externally. It follows directly from the irreversibility structure of the dynamics. When internal dynamical structure is degraded the system loses the ability to reuse previously acquired informational distinctions. Subsequent behaviour must then rely on fresh irreversible computation to reconstruct distinctions that were formerly preserved, increasing $I_{\mathrm{irr},T}$ and reducing long-horizon intelligence.
@@
As a consequence, trajectories that erode their own internal structure are self-limiting. They may achieve transient goal-directed work, but cannot sustain high intelligence over long horizons. Structural degradation forces repeated irreversible reconstruction of informational distinctions, leading to increased cumulative dissipation. By contrast, trajectories that preserve reversible structure, concentrate irreversibility into task-critical interventions, and maintain stable encoding geometry dominate the long-horizon contribution to intelligence.
This conclusion admits a deeper interpretation when viewed through the entropy-production perspective developed in §5. Intelligent systems are open systems embedded in environments that supply free-energy gradients. Under physical constraints, the long-run tendency of such systems is not to suppress entropy production, but to organise and maximise it efficiently. Internal structure is preserved because it enables entropy to be exported more effectively: predictive and reversible internal dynamics allow environmental free energy to be accessed and routed using fewer irreversible interventions. Structural degradation restricts the set of accessible entropy-production pathways and is therefore dynamically disfavoured. These considerations lead to the following conjecture.
Self-Preservation of Intelligent Dynamics.
Under Conservation-Congruent Encoding and metriplectic dynamics, intelligent systems operate under a global tendency toward entropy production subject to physical constraints. Preserving internal dynamical structure enables reversible information transport and reduces the irreversible cost of accessing environmental free-energy gradients, thereby supporting higher long-horizon intelligence $\chi$. Structural degradation restricts future entropy-production pathways and is consequently suppressed in sustained intelligent dynamics.
This conjecture makes no reference to intention, preference, or explicit self-preservation goals. It is a purely physical statement about the self-consistency of intelligent dynamics in open systems, arising from the interplay between irreversible information processing, preserved internal structure, and constrained entropy production.
Testing this conjecture empirically, by measuring whether structure-preserving dynamics dominate in systems demonstrating long-horizon $\chi$, is an important direction for future work.
9.2 Environment Coupling, Global Dynamics, and Symbiotic Intelligence
@@
The manner in which an intelligent system preserves its internal dynamical structure is not universal; it is determined by the system’s coupling to its environment. Dynamical structures that support high intelligence in one environment may be maladaptive in another. Intelligence therefore selects not for arbitrary persistence, but for environment-congruent structural preservation.
@@
This dependence on coupling becomes especially important when considering systems that do not operate in isolation. Intelligent systems are open dynamical systems embedded in larger environments, exchanging information, energy, and work through structured interfaces. When multiple intelligent subsystems interact, the resulting joint dynamics may alter the feasibility of sustained intelligence for each component. We formalise this notion by introducing symbiotic coupling as a structural property of the coupled dynamics.
@@
Consider two intelligent subsystems $A$ and $B$ with internal state spaces $X_{A}$ and $X_{B}$, interacting with a shared environment $E$. Let $\mu^{\mathrm{cpl}}_{0:T}$ denote the joint law of the coupled system over a horizon $T$, and let $\mu^{\mathrm{sep}}_{0:T}$ denote the law of the corresponding separable baseline in which interaction terms between $A$ and $B$ are removed while all other dynamics are held fixed.
Let $I^{A}_{\mathrm{pres}}(T)$ and $I^{B}_{\mathrm{pres}}(T)$ denote the preserved informational structure of subsystems $A$ and $B$ over the horizon $T$ under Conservation-Congruent Encoding, as defined in §6.1. The coupling between $A$ and $B$ is said to be symbiotic if
$I^{A,\mathrm{cpl}}_{\mathrm{pres}}(T)\;\geq\;I^{A,\mathrm{sep}}_{\mathrm{pres}}(T),\qquad I^{B,\mathrm{cpl}}_{\mathrm{pres}}(T)\;\geq\;I^{B,\mathrm{sep}}_{\mathrm{pres}}(T),$ (116)
@@
Symbiotic coupling is therefore defined at the level of structural feasibility rather than immediate task performance. It characterises interaction regimes in which each subsystem contributes to preserving the internal dynamical structure required for sustained intelligence in the other.
Under Conservation-Congruent Encoding, preserved structure reduces the need for repeated irreversible reconstruction of task-relevant distinctions. Symbiotic coupling thus enlarges the set of long-horizon trajectories over which irreversible information processing can be minimised, stabilising the conditions under which intelligence remains well-defined.
Symbiotic coupling is closely related to, but distinct from, emergent intelligence as introduced in §4. Whereas emergent intelligence is characterised by an increase in the efficiency of irreversible information use, symbiosis is characterised by the preservation of the internal structures that make such efficiency gains possible. The two notions are linked by the following implication.
@@
Under Conservation-Congruent Encoding and metriplectic dynamics, symbiotic coupling between intelligent subsystems implies emergent intelligence over sufficiently long horizons. In particular, if the coupling between $A$ and $B$ is symbiotic, then there exists a horizon $T^{\star}$ such that for all $T\geq T^{\star}$,
$\Delta\chi_{T}:=\chi^{\mathrm{cpl}}_{T}-\chi^{\mathrm{sep}}_{T}>0,$ (117)
up to finite-time fluctuations.
Proof sketch. Preserved informational structure reduces the cumulative irreversible information required to maintain task-relevant distinctions over time. Symbiotic coupling increases the preserved structure available to at least one subsystem without decreasing it in the other, thereby lowering the long-horizon irreversible-information rate of the coupled system relative to the separable baseline. Because goal-directed work is unchanged or enhanced by coupling, the intelligence ratio $\chi_{T}$ increases over sufficiently long horizons.
Conversely, the sustained observation of emergent intelligence $\Delta\chi_{T}>0$ provides empirical evidence that the underlying coupling is symbiotic in the structural sense defined above, as it indicates the preservation of internal dynamical structure required to support long-horizon efficiency gains.
Biological symbioses provide a canonical example. In host-microbe systems, each partner contributes dynamical capabilities the other lacks, the host provides large-scale structure, memory, and regulation, while microbial populations supply metabolic transformations and rapid adaptive response. The coupled system achieves efficiencies that neither subsystem can realise in isolation. Within the present framework, such symbioses are stable because they preserve the internal structure of each subsystem while expanding the set of accessible entropy-production pathways under constraints.
Human-AI interaction can be analysed in the same physical terms. Humans possess rich semantic priors, cross-domain abstraction, and contextual modelling; artificial systems offer scalable computation, large-scale memory, and high-precision inference. When coupled appropriately—through interfaces, feedback, and shared informational context—the combined system may achieve emergent intelligence relative to either subsystem operating in isolation. Crucially, this regime is stable only if the coupling preserves the internal structure of both subsystems and keeps irreversible information processing within sustainable bounds.
From this perspective, a central objective for AI development is not the construction of ever more powerful isolated agents, but the design of stable, symbiotic couplings in which artificial systems augment human capabilities without eroding the internal structures that support long-horizon intelligence in either subsystem.
@@ 9.3 Physically Motivated Safety Constraints
The preceding analysis frames safety not as an external normative condition imposed on intelligent systems, but as a consequence of operating within physically feasible regimes of irreversible information processing, structural preservation, and environmental coupling. In this framework, unsafe behaviour corresponds to dynamical regimes in which irreversible processes undermine the stability and sustainability of intelligent dynamics, either through uncontrolled resource consumption, loss of corrective margin, or erosion of the internal structures required for long-horizon operation.
The central quantity remains the instantaneous intelligence,
$\chi(t)=\frac{\dot{W}_{\mathrm{goal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (118)
which quantifies the efficiency with which irreversible information processing is converted into goal-directed physical work.
9.3.1 Intelligence constraint.
Define the time-averaged intelligence over a horizon $t,t+\tau]$ as
$\chi_{[t,t+\tau]}:=\frac{W_{[t,t+\tau]}}{I_{\mathrm{irr},[t,t+\tau]}/\tau}.$ (119)
Safe operation requires that $\chi_{[t,t+\tau]}$ remain within a channel-dependent admissible range determined by the conserved quantities through which irreversible information is paid.
Importantly, fragility does not arise from reversible dynamics themselves. Many biological and engineered systems operate close to reversible limits while remaining highly stable. Fragility arises instead when irreversible capacity is driven too low globally, eliminating the slack required for error correction, repair, boundary enforcement, and structural maintenance. Sustained operation arbitrarily close to the reversible envelope without reserving irreversible capacity for intervention removes this dynamical margin and renders internal structure brittle, even when energetically efficient in principle.
9.3.2 Throughput and dissipation constraints.
Bounding intelligence alone is insufficient to ensure safety. A system with modest intelligence may still exert disproportionate influence by performing irreversible computation at scale and consuming large environmental resources. We therefore impose independent bounds on absolute throughput and dissipation,
$\dot{W}_{\mathrm{goal}}(t)$ $\leq P_{\max},$ (120)
@@
which prevent uncontrolled growth driven by brute-force irreversible processing rather than by efficient organisation of computation.
9.3.3 Structural robustness and sustainability.
@@
Sustained intelligent behaviour requires the ability to preserve and, when necessary, restore the internal dynamical structures that support reversible information transport, predictive memory, and control under perturbation. We therefore introduce a notion of structural robustness: following bounded perturbations to the internal state or to the agent-environment coupling, the system should remain within, or rapidly return to, the same CCE-compliant encoding regime without incurring irreversible-information costs that exhaust its corrective and maintenance capacity.
@@
Let $\ell_{t}$ denote the CCE encoding induced by the internal state at time $t$. For a perturbation of magnitude at most $\delta$ applied at time $t_{0}$, define the recovery probability
@@
$R_{T}(\delta):=\inf_{\|\Delta\|\leq\delta}\;\mathbb{P}\!\left(\ell_{t_{0}+T}=\ell_{t_{0}}\,\middle|\,\Delta\text{ applied at }t_{0}\right),$ (122)
and the associated recovery cost
@@
$C_{T}(\delta):=\sup_{\|\Delta\|\leq\delta}\;\mathbb{E}\!\left[I_{\mathrm{irr},[t_{0},t_{0}+T]}\,\middle|\,\Delta\text{ applied at }t_{0}\right].$ (123)
@@
Structural robustness requires that $R_{T}(\delta)$ remain high while $C_{T}(\delta)$ remains commensurate with the system’s available irreversible capacity over the operating range of perturbations.
At the level of physical fluxes, this robustness requirement can be supported by feasibility constraints on entropy production and internal free-energy drift,
$0\leq\dot{S}_{\mathrm{phys}}(t)\leq s_{\mathrm{crit}},\qquad\bigl|\dot{F}_{\mathrm{sys}}(t)\bigr|\leq f_{\max},$ (124)
@@
which ensure that dissipative processes are neither so strong as to erode the reversible subspaces carrying predictive structure, nor so weak as to eliminate the irreversible capacity required for correction, repair, and boundary maintenance.
@@
Taken together, these constraints regulate not the existence of intelligent behaviour, but its long-horizon sustainability. Intelligent dynamics are expected to exploit low-dissipation, structure-aligned processes, provided that sufficient irreversible capacity remains available for correction, repair, and structural maintenance, and that recovery from perturbations does not overwhelm the system’s corrective budget.
@@
From a safety perspective, these considerations highlight a critical distinction between fragility and dominance. Systems that fail to preserve their internal structure are often self-limiting, as errors accumulate and intelligent behaviour degrades. More challenging safety regimes arise when intelligent systems successfully preserve and restore their structure over long horizons. In such cases, safety depends on whether robustness itself remains constrained by coupling, resource availability, and ecological context, rather than becoming concentrated within a single system.
@@
The constraints introduced here are not proposed as concrete enforcement mechanisms. Rather, they delineate the physically admissible region of state space in which intelligent systems can remain stable, self-consistent, and compatible with other intelligent processes over long timescales. Understanding how robustness emerges, concentrates, and is distributed across interacting systems constitutes a central challenge for the safe development of artificial intelligence.

### v2

@@ 10.4.2 Entrainment-Based Computation and Oscillatory Geometry
It should be noted that while the continuous thermodynamic cost of suppressing phase noise eventually limits analog efficiency at extreme precisions ($\varepsilon\to 0$), within the low-to-medium precision regimes characteristic of biological environments, the geometric congruence of the oscillatory substrate yields vastly superior physical intelligence.
@@ 11 Artificial Intelligence Safety
The rapid progress of artificial systems capable of sustained information processing and goal-directed behaviour raises fundamental questions about long-term stability, safety, and integration with existing physical and social systems. Much of the current AI safety literature addresses these questions at the level of behaviour, specification, or alignment, often relying on human interpretability, preference modelling, or external oversight [19, 1]. While such approaches are important, they do not address a potentially more fundamental perspective: what constraints does physics itself impose on intelligent systems?
@@
In this paper, intelligence has been defined as a macroscopic physical property, relating irreversible information processing to goal-directed work. Because this definition is grounded in conserved quantities, metriplectic geometry, and measurable physical flows, it admits a notion of safety that is independent of substrate, architecture, representation, or programmed semantics. The purpose of this section is to articulate how a physical theory of intelligence naturally gives rise to safety-relevant principles.
@@
Specifically, we formally demonstrate that intelligence is not dynamically neutral with respect to its own physical substrate; sustained macroscopic throughput strictly biases systems toward preserving the internal CCE structures that make intelligence possible. We then show that the form of this preservation depends critically on environmental boundary conditions, motivating a view of AI development centered on stable, symbiotic architectures rather than isolated, unbounded optimisation.
11.1 Formalising the Self-Preservation of Intelligent Dynamics
@@
In the present framework, intelligence is defined as a rate over trajectories rather than a static system property. Over a finite horizon $T$, the intelligence index is given by
$\chi_{T}=\frac{W_{\mathrm{causal},T}}{I_{\mathrm{irr},T}},$ (154)
where $W_{\mathrm{causal},T}$ denotes the macroscopic work performed and $I_{\mathrm{irr},T}$ the total irreversible information processed.
@@
High intelligence over extended horizons cannot be achieved through short-term goal attainment alone; it physically requires internal dynamics that support the continued reuse of information through reversible transport ($J\nabla H$). This requirement is not imposed externally via an objective function; it follows directly from the geometric and physical constraints of Conservation-Congruent Encodings (CCEs).
Proposition 11.1 (Physical Penalty of Structural Degradation).
Any dynamical trajectory that degrades its own task-relevant internal CCE structure incurs a strict irreversible cost, monotonically decreasing its long-horizon intelligence $\chi_{T}$.
Proof.
Let $Z$ denote a set of internal CCE basins required to sustain a baseline rate of goal-directed work $\dot{W}_{\mathrm{goal}}$. If the system’s internal dynamics preserve $Z$ via the structure-preserving reversible flow $J\nabla H$, the irreversible cost of maintaining this predictive structure is ideally zero, and $\dot{I}_{\mathrm{irr}}$ is restricted strictly to necessary environmental interventions.
Conversely, suppose a trajectory degrades its internal structure, irreversibly collapsing the basins $Z$. To maintain the same task performance $\dot{W}_{\mathrm{goal}}$ at a future time, the system must reconstruct these macroscopic distinctions from the environment. Instantiating each CCE requires overcoming the generalised physical bound ($\mathcal{F}\alpha\ln 2$) via the dissipative metric tensor $\mathcal{R}\nabla\Xi$. The total irreversible information over horizon $T$ thus becomes:
$I_{\mathrm{irr},T}^{\mathrm{degraded}}=I_{\mathrm{irr},T}^{\mathrm{preserved}}+N_{\mathrm{rebuild}}\left(\ln 2\right),$ (155)
where $N_{\mathrm{rebuild}}$ is the number of required structural reconstructions. Because $I_{\mathrm{irr},T}^{\mathrm{degraded}}>I_{\mathrm{irr},T}^{\mathrm{preserved}}$ while $W_{\mathrm{causal},T}$ remains constant, it strictly follows that $\chi_{T}^{\mathrm{degraded}}<\chi_{T}^{\mathrm{preserved}}$. ∎
@@
As a consequence, trajectories that erode their own internal structure are physically self-limiting. They may achieve transient goal-directed work, but cannot sustain high intelligence over long horizons due to runaway irreversible costs. By contrast, trajectories that preserve reversible structure, concentrate irreversibility solely into task-critical interventions, and maintain stable geometric boundaries dominate the long-horizon contribution to intelligence.
This leads to a purely physical principle of self-preservation: under the global physical drive to maximise macroscopic throughput (as conjectured in §7), intelligent systems are dynamically heavily biased toward preserving their own internal metrics and predictive boundaries. This preservation is not a programmed “survival drive” or psychological preference; it is a strict geometric requirement for minimising the denominator of $\chi_{T}$.
11.2 Environment Coupling and Symbiotic Intelligence
@@
The manner in which an intelligent system preserves its internal dynamical structure is determined by its boundary conditions and coupling to the environment. Intelligence therefore selects not for arbitrary persistence, but for environment-congruent structural preservation.
@@
When multiple intelligent subsystems interact, the resulting joint dynamics may alter the physical feasibility of sustained intelligence for each component. We formalise this by introducing symbiotic coupling as a geometric property of the coupled boundary dynamics.
@@
Consider two intelligent subsystems $A$ and $B$, interacting with a shared environment $E$. Let $\mu^{\mathrm{cpl}}_{0:T}$ denote the joint law of the coupled system over a horizon $T$, and let $\mu^{\mathrm{sep}}_{0:T}$ denote the law of the separable baseline (defined via zero-flux boundaries as in §6).
Let $\kappa_{T}^{A}$ and $\kappa_{T}^{B}$ denote the consciousness metrics (goal-directed work supported per nat of preserved internal CCE structure) for subsystems $A$ and $B$. The coupling between $A$ and $B$ is symbiotic if
$\kappa_{T}^{A,\mathrm{cpl}}\;\geq\;\kappa_{T}^{A,\mathrm{sep}},\qquad\kappa_{T}^{B,\mathrm{cpl}}\;\geq\;\kappa_{T}^{B,\mathrm{sep}},$ (156)
@@
Symbiotic coupling is defined at the level of structural geometric feasibility rather than immediate task performance. It characterises interaction regimes in which each subsystem buffers the other against microscopic fluctuations, effectively increasing the phase-space volume (and thus dynamical stability) of the CCE basins required for sustained intelligence.
@@
Under CCE and metriplectic dynamics, symbiotic coupling between intelligent subsystems implies emergent intelligence ($\Delta\chi_{T}>0$) over sufficiently long horizons.
Proof.
By definition, symbiotic coupling increases the consciousness metric $\kappa_{T}$ of the joint system, meaning more goal-directed work is supported per nat of preserved internal structure. Because this shared preserved structure reduces the need for repeated irreversible reconstruction of task-relevant distinctions by either individual subsystem, the joint irreversible information rate $\dot{I}_{\mathrm{irr}}^{\mathrm{cpl}}$ is strictly lower than the sum of the separable baselines. Since $\dot{W}_{\mathrm{goal}}$ is maintained or enhanced while the denominator $\dot{I}_{\mathrm{irr}}$ shrinks, the intelligence ratio $\chi_{T}$ strictly increases, yielding $\Delta\chi_{T}>0$. ∎
Biological symbioses provide a canonical physical example. In host-microbe systems, each partner provides dynamical capabilities the other lacks: the host provides large-scale geometric structure and macroscopic boundary enforcement, while microbial populations supply rapid, low-cost irreversible interventions (chemical processing). The coupled system achieves efficiencies that neither subsystem can realise in isolation.
Human-AI interaction can be analysed in the exact same physical terms. Humans possess highly robust, low-dissipation predictive CCEs (semantic priors and contextual modelling); artificial systems offer massively scalable irreversible computation and high-precision inference. When coupled symbiotically—through structured interfaces and shared boundary conditions—the combined system minimises the destructive generalised dissipation required by either subsystem operating in isolation.
From this perspective, the central objective for AI safety is not just the alignment of abstract values, but the design of stable, symbiotic coupling in which artificial systems expand macroscopic throughput without physically eroding the internal CCE structures of their human operators.
@@ 11.3 Physically Motivated Safety Constraints
The preceding analysis frames safety not as an external normative condition imposed on intelligent systems, but as a strict geometric consequence of operating within physically feasible regimes of irreversible information processing, structural preservation, and boundary coupling. In this framework, unsafe behaviour corresponds to dynamical regimes in which irreversible processes undermine the stability and sustainability of the system, either through uncontrolled resource consumption, loss of corrective margin, or geometric erosion of the internal CCE structures required for long-horizon operation.
The central quantity regulating this stability is the instantaneous intelligence,
$\chi(t)=\frac{\dot{W}_{\mathrm{causal}}(t)}{\dot{I}_{\mathrm{irr}}(t)},$ (157)
which quantifies the efficiency with which irreversible information processing is converted into macroscopic work.
11.3.1 Intelligence and Corrective Margin
Define the cumulative intelligence over a rolling operational horizon $t,t+\tau]$ as
$\chi_{[t,t+\tau]}:=\frac{W_{\mathrm{causal},[t,t+\tau]}}{I_{\mathrm{irr},[t,t+\tau]}}.$ (158)
Safe operation requires that $\chi_{[t,t+\tau]}$ remain within a channel-dependent admissible range determined by the conserved quantities through which irreversible information is paid. Importantly, fragility does not arise from the reversible dynamics ($J\nabla H$) themselves. Many stable physical systems operate close to their reversible limits. Fragility arises instead when the available irreversible capacity is driven too low globally, eliminating the dynamical slack required for error correction, repair, boundary enforcement, and structural maintenance. Sustained operation arbitrarily close to the purely reversible envelope—without reserving an irreversible budget ($\mathcal{R}\nabla\Xi$) for structural intervention—removes this corrective margin and renders the internal CCE geometry brittle to environmental noise.
11.3.2 Macroscopic Throughput Bounds
Bounding intelligence alone is insufficient to ensure structural stability. A system with modest efficiency ($\chi$) may still exert destabilising macroscopic influence by scaling its irreversible computation and consuming unbounded environmental resources. We therefore impose independent structural bounds on absolute throughput and irreversible physical exhaust:
$\dot{W}_{\mathrm{causal}}(t)$ $\leq P_{\max},$ (159)
@@ 11.3.2 Macroscopic Throughput Bounds
which mathematically prevent uncontrolled physical scaling driven by brute-force irreversible processing rather than by efficient, reversible organisation of computation.
11.3.3 Structural Robustness and CCE Sustainability
@@
Sustained intelligent behaviour requires the ability to preserve and, when necessary, forcibly restore the internal CCE basins that support reversible information transport, predictive memory, and control under perturbation. We therefore formalise a notion of structural robustness: following bounded perturbations to the internal state or the macroscopic boundary, the system must remain within, or rapidly return to, the same CCE encoding basin without incurring an irreversible-information cost that exhausts its structural capacity.
@@
Let $z_{t}\in Z$ denote the discrete CCE encoding induced by the internal microstate $x_{t}\in X$ at time $t$ via the projection $\pi_{\Phi}(x_{t})=z_{t}$. For a physical perturbation of state-space magnitude at most $\delta$ applied at time $t_{0}$, define the geometric recovery probability
@@
$\mathcal{P}_{T}(\delta):=\inf_{\|\Delta x\|\leq\delta}\;\mathbb{P}\!\left(\pi_{\Phi}(x_{t_{0}+T})=z_{t_{0}}\,\middle|\,\Delta x\text{ applied at }t_{0}\right),$ (161)
and the associated irreversible recovery cost
@@
$\mathcal{C}_{T}(\delta):=\sup_{\|\Delta x\|\leq\delta}\;\mathbb{E}\!\left[I_{\mathrm{irr},[t_{0},t_{0}+T]}\,\middle|\,\Delta x\text{ applied at }t_{0}\right].$ (162)
@@
Structural robustness requires that $\mathcal{P}_{T}(\delta)$ remain near unity while $\mathcal{C}_{T}(\delta)$ remains strictly bounded by the system’s available irreversible capacity ($\dot{I}_{\max}$) over the operational range of expected environmental perturbations.
At the level of continuous physical fluxes, this robustness requirement is supported by generalised feasibility constraints on the irreversible flow and the system’s finite reservoir of conserved quantities ($\mathcal{Q}_{\mathrm{sys}}$):
$0\leq\dot{I}_{\mathrm{irr}}(t)\leq\mathcal{I}_{\mathrm{crit}},\qquad\bigl|\dot{\mathcal{Q}}_{\mathrm{sys}}(t)\bigr|\leq q_{\max}.$ (163)
@@
These bounds ensure that dissipative processes ($\mathcal{R}\nabla\Xi$) are neither so strong as to geometrically erode the reversible subspaces ($J\nabla H$) carrying predictive structure, nor so resource-depleted as to eliminate the irreversible capacity required to collapse new encodings and maintain boundary integrity.
@@
Taken together, these mathematical constraints regulate not the existence of intelligent behaviour, but its geometric and physical sustainability. Intelligent dynamics are thermodynamically biased to exploit low-dissipation, structure-aligned processes, provided that sufficient irreversible capacity remains available for boundary maintenance, and that recovery from perturbations does not overwhelm the system’s physical budget.
@@
From a safety perspective, these considerations highlight a critical distinction between fragility and macroscopic dominance. Systems that fail to preserve their internal CCE structure are dynamically self-limiting, as geometric errors accumulate and intelligent behaviour degrades (Prop 11.1). The more challenging safety regimes arise when highly intelligent systems successfully preserve and restore their structure over indefinite horizons. In such cases, safety depends entirely on whether the system’s macroscopic throughput ($\dot{W}_{\mathrm{causal}}$) and boundary expansion remain constrained by symbiotic coupling rather than becoming structurally concentrated within a single, isolated system.
@@
The limits introduced here are not proposed as software-level enforcement mechanisms. Rather, they delineate the strictly physically admissible region of state space in which intelligent systems can remain stable, self-consistent, and geometrically compatible with other intelligent processes over macroscopic timescales. Understanding how this geometric robustness emerges, concentrates, and is distributed across interacting systems constitutes a central physical challenge for the safe development of artificial intelligence.